Methane gas solubility
a. If methane (CH4) makes up 0.000181% of dry air at 25C, what is the concentration of methane in water in molarity? Assume the partial pressure of water is 0.0313atm at 25C.
b. Typically, small concentrations are reported not in molarity but parts per million (ppm). PPM is equal to mg/L in water. Based on that, what is the concentration of methane from above in ppm?
Answer #1

Molarity = Moles of solute/L of solvent

Lets assume that we have 1L of air at 25C

Methane is in very small proportion compared to water, its contribution to total pressure can be ignored.

a. At 25C, 22.4L of air has 1 mole of air(Avogadro's Law) , so 1L air will have 1/22.4 = 0.04464 Moles of air , partial Pressure of water is 0.0313atm , so there are 0.0313*0.04464 moles of water. =  0.001397 mol water vapor in the gaseous phase ,

Moles of Methane per L of water vapor= 0.00000181 * 0.04464 = 8.08*10-7 Mol/L

b. Lets calculate mg/L further from our previous answer. Methane atomic mass can be assumed to be 16.

mg of methane = 1000*g of methane = 1000* 8.08*10-7*16= 0.129 ppm